OpenStudy (anonymous):

If 5x -y= x +36 and y= 2x, x=

OpenStudy (anonymous):

5x -y= x +36 y= 2x We know what y is, so we can go ahead and fit that into the first equation. 5x- 2x=x+36 Subtract the two like terms 3x=x+36 -x -x Subtract x from both sides 2x=36 /2 /2 Divide by 2 on both sides x=18